# Channel Measurement, Modeling, and Simulation for 6G: A Survey and Tutorial

### v3

@@ VI RIS Channel Measurement and Modeling
Different from traditional communication technologies, an RIS realizes active control of a wireless signal propagation environment by passively reflecting electromagnetic waves. Fig. 17 shows a typical RIS-assisted communication scenario. In this scenario, the LOS path is obstructed, and an RIS is employed to enhance the received signal. Obviously, traditional channel modeling methods are not suitable for this kind of channel, which requires new research on the RIS channel model. Notably, in Fig. 17, the RIS channel was characterized by a Tx-RIS-Rx cascaded channel, which was composed of a Tx-RIS subchannel, RIS-Rx subchannel, and RIS response. In this cascaded channel, the subchannels required independent modeling and were combined in the RIS response. The subsequent RIS channel modeling described in this paper is based on this method.
@@ VI-A RIS Channel Measurements
Some of the other measurement campaigns are carried out in practical scenarios. A near-field RIS channel measurement is carried out in indoor and outdoor environments to research the path loss characteristics of RIS channel [208]. Furthermore, to facilitate channel modeling in system simulations, research [209] proposes a reflective RIS physical model and conducts indoor RIS measurement. This work aims to abstract the RIS physical model for application in practical channel modeling scenarios. In [210], a RIS-assisted channel measurement platform that can be applied to various complex scenarios based on the time domain correlation principle is built. Using this platform, the authors carry out a channel measurement in a typical indoor corridor scenario at 26 GHz, as shown in Fig. 18. They research the path loss characteristic of RIS-assisted channels in corridor scenarios based on the measurement data.
@@
Discrete 5.8 Indoor, outdoor Power gain [208]
Discrete 5.8
@@
Discrete 26 Indoor PL [210]
Discrete 26.9 Indoor Radiation pattern [209]
Continuous 4.25, 10.5 Microwave anechoic chamber Power gain [211]
Continuous 27, 33 Microwave anechoic chamber
Power gain,
radiation pattern
[206]
@@ VI-B RIS Channel New Characteristics
In the RIS channel, there are three types of characteristics that are worth paying attention to, namely the reflection coefficient characteristics of RIS element, RIS radiation pattern characteristics, and RIS cascade loss characteristics. Among them, RIS radiation pattern characteristics consists of RIS overall radiation pattern characteristics and RIS element radiation pattern characteristics. The reasons are as follows.
Firstly, the RIS is one of the most important nodes in the RIS channel, and its response are an important component of the RIS channel. In order to calculate the RIS response, it is necessary to calculate the reflection coefficient of each RIS element in advance. Thus, the RIS element reflection coefficient characteristics are very important. Secondly, the final form of RIS response is RIS overall radiation pattern or RIS elemenmt radiation pattern, therefore, their characteristics can not be ignored either. Fig. 19 shows a feasible process of calculating the RIS radiation pattern, corresponding to (18)-(23). Finally, since the RIS cascaded channel is obtained by cascading Tx-RIS sub-channel, RIS response, and RIS-Rx sub-channel, its loss characteristics are different from traditional communication channel. Thus, RIS cascade loss characteristics are worth studying.
@@ VI-B2 RIS radiation pattern characteristics
VI-B3 RIS cascade loss characteristics
Multiple studies have been conducted on the RIS cascade loss of the RIS channel based on theoretically derived and measurements. They provide rich RIS cascade loss modeling methods and combined reveal the RIS cascade loss characteristics of the RIS channel.
During a period, the relationship between the RIS cascade loss of the RIS channel and the distances $d_{1}$ and $d_{2}$ is very relevant, where $d_{1}$ is the distance between the transmitter and RIS, and $d_{2}$ is the distance between the RIS and receiver. A theoretical RIS cascade loss model in free space under the far-field condition was proposed in [218]. The theoretical analysis demonstrated that the RIS cascade loss is proportional to $(d_{1}d_{2})^{2}$. Similar conclusions were also obtained in measurements of a corridor scene. Authors [210] carried out a channel measurement in a typical indoor corridor scenario and studied the RIS cascade loss characteristic of the RIS channel in corridor scenarios. The results indicate that a modified version of the multiplicative distance model can effectively fit the measurement data.
Furthermore, some different conclusions can be made. In [206], a general RIS cascade loss model and some special models in free space were proposed. These special models apply to RIS-assisted far-field beamforming, RIS-assisted near-field focusing, and RIS-assisted near-field broadcasting cases. The RIS-assisted far-field beamforming and RIS-assisted near-field broadcasting models were verified by several measurements conducted in a microwave chamber at 27 and 33 GHz. Different from above, the authors found that the RIS cascade loss is proportional to $(d_{1}+d_{2})^{2}$ in the RIS-assisted near-field broadcasting case. This indicates that the RIS cascade loss characteristics of the RIS channel are related to the encoding method of the RIS and depend on whether the terminal is in the far field or near field.
@@ VI-C RIS Channel Modeling
Considering the prevalent use of reflective RIS, it is often placed between the transmitter and receiver, resulting in a cascaded channel known as the Tx-RIS-Rx channel. The main objective of this subsection is to construct an effective cascaded RIS channel model to accurately capture the response characteristics of the RIS and channel properties, while balancing the model complexity and accuracy.
@@
In [205], the authors propose a GBSM-based channel modeling framework for the RIS cascaded channel, which takes into account the incident angle dependence of the RIS element reflection coefficient. The model was applied to an RIS-assisted MIMO communication scenario, as shown in Fig. 17. The CIR of the RIS cascaded channel between the $s^{\textrm{th}}$ transmitter and the $u^{\textrm{th}}$ receiver is expressed as (24). Similar to the channel model presented in (6), most of the channel parameters in the RIS cascade channel model are defined similarly. However, the RIS cascaded channel model differs in that it includes two subchannels and the RIS response. The specific parameters of the RIS cascaded channel model are listed below.
@@
Implementation of the proposed model is based on the 3GPP standardized channel model with some improvements, including three parts: 1) general and large-scale parameters, 2) small-scale parameters, and 3) the CIR for RIS cascaded channel. The model generates large-scale parameters (LSPs) and SSPs for both the Tx-RIS and RIS-Rx subchannels. The overall radiation pattern of the RIS is calculated using a method proposed in [216] to account for the incident angle-dependent characteristics of the RIS reflection coefficient. For further details on the specific implementation processes, readers may refer to [205].
@@ VI-D Summary and Prospects
The RIS channel measurement is primarily aimed at measuring the power gain caused by RIS. The characteristics of the RIS channel primarily include reflection coefficient characteristics of the RIS element, RIS radiation pattern characteristics, and RIS cascade loss characteristics. These characteristics are worth considering in RIS channel modeling. Based on these findings, we foresee some challenges that RIS channel modeling research could encounter.
@@ VI-D1 RIS channel measurement
RIS channel measurement plays a crucial role in understanding RIS channel characteristics and validating RIS channel models. However, due to the high transmission loss of the RIS-assisted channel caused by the Tx-RIS-Rx-cascaded link, it imposes greater requirements on the transmission power and dynamic range of the measurement platform. In addition, the measurement platform needs a complete architecture, such as multiple channels, a high sampling rate, and transmit-receive separation, to capture channel characteristics in various complex scenarios.
@@ VI-D3 RIS channel modeling
The passive characteristic makes RIS have the advantage of low cost and low power consumption. But this also leads to a problem where RIS requires a large area to provide sufficient gain. Therefore, the RIS channel may exhibit near-field effect and SnS like E-MIMO channel. In this case, the RIS response can not be characterized by the RIS overall radiation pattern. Thus, modeling near-field effects and SnS in the RIS channel is a major challenge.
@@ VII SAGIN Channel Measurements and Models
One of the 6G usage scenarios is ubiquitous connectivity by addressing the challenges of connectivity, coverage, capacity, data rate and the mobility of terminals [6]. One focus of this usage scenario is to address presently uncovered or scarcely covered areas, and SAGIN is considered a resolution to achieve worldwide coverage. SAGIN is an integration of satellite systems, aerial networks, and terrestrial communications. Thus in SAGIN, channel characteristics in the communication scenarios of different segments (space, air and ground) should be considered. However, due to the long propagation distance and high node mobility, SAGIN channel characteristics are distinctly different from those of the traditional terrestrial channels [246]. Thus in this section, we review and introduce the present research on channel measurements, characteristics, and models of SAGIN. Since terrestrial systems have been discussed in the above sections, this section is focused more on the SG and AG scenarios. In addition, a brief summary and potential future research directions are provided.
@@ VII-A SAGIN Channel Measurements
For clarity, we classified SAGIN channel measurements, into SG, AG, and air-to-air (AA), and summarized them in Table V.
@@
For SG channels, research work has been started in the early years. Earlier, researchers [223] conducted measurements using the ATS-6 satellite to study the propagation loss of SG links at 860 and 1550 MHz. In [224], propagation loss measurements were carried out at several geographical ground locations in the Q/V frequency band, and a low earth orbit (LEO) satellite channel model was developed using machine learning and statistical methods. A series of measurements of tree attenuation were conducted and studied in various frequency bands such as L-band and K-band [225, 226, 227, 228, 229, 230, 231]. In [232], a satellite MIMO channel study was conducted to investigate the channel capacity in an urban scenario. In [235], meteorological effects, such as rain, clouds, and scintillation, were studied based on field channel measurements in a tropical area. In addition, researchers in [233, 234] conducted measurements to study rain attenuation at the Ka/Q band. According to Table V, it can be found that most of the channel measurements of SG were carried out at L-band, S-band, K-band and Q-band [247], which are the common frequency bands for satellite communications. For channel characteristics, these measurement campaigns mainly focus on the path loss, shadow fading (especially shadow fading caused by tree attenuation), and meteorological effects, and less attention is paid to the small scale characteristics of the channel.
@@
For aerial channels like AG and AA, due to the rise of UAV technology, the corresponding channel measurements have gradually conducted. In [241], the path loss of the UAV channel was studied at 2.3 GHz in urban and nonurban scenarios. In [237, 238, 239], the National Aeronautics and Space Administration (NASA) Glenn Research Center sponsored a comprehensive AG channel measurement study of the C-bands. These measurements covered nearly all the typical local ground environments, including overwater, hilly, mountainous, suburban, and near-urban areas. Researchers in [240] also conducted wideband measurements using a direct-sequence spread-spectrum channel measurement platform to characterize propagation at 2.05 GHz. In [236], a channel measurement for a low-altitude AG channel was conducted at 915 MHz. In [242], AG channel measurements were conducted in open and suburban areas at 4.3 GHz. In [244], authors studied time-varying AG channels based on the measurement data at 6.5 GHz, with a particular focus on the clustering and tracking of multipath components. In [243], the measurement of the AG channel in a near-airport scenario at 5 GHz was conducted. In addition, AA channels for UAVs were investigated based on laboratory measurements and several experimental flights in [245]. According to Table V, it can be found that measurements of aerial channels are conducted mainly in AG scenarios, channel measurements in AA and other scenarios are scarce. Besides, measurement campaigns were conducted in relatively low frequency bands, which are from 0.92 to 6.5 GHz. For channel characteristics, in addition to path loss and shadow fading, AG channel measurements also focus on small scale characteristics such as delay spread, Ricean K-factor (K-factor), Doppler spread and others.
@@ VII-C SAGIN Channel Modeling
where $r$ is the received signal envelope, $ze^{j\phi_{0}}$ represents the LOS component, $we^{j\phi}$ represents other multipaths, the phases $\phi_{0}$ and $\phi$ are uniformly distributed between 00 and $2\pi$, $z$ is log-normally distributed , and $w$ has a Rayleigh distribution. In recent years, meteorological effects have been gradually taken into consideration. In [253], a channel model taking into account weather impairments was proposed. In this case, the PDF of the received signal amplitude is
@@
$K_{\rm total}=K_{\rm LOS}-K_{\rm Rain}-K_{\rm Cloud}$ (27)
@@ VII-D3 SAGIN channel modeling
Channel simulator is the software to generate CIR in different scenarios and configurations based on the existing modeling framework and configuration parameters. Channel simulator plays an important role in designing, evaluating, and optimizing communication systems. In this section, a 6G channel simulator named Beijing University of Posts and Telecommunication and China Mobile Communications Corporation Channel Model Generator-IMT2030 (BUPTCMCCCMG-IMT2030) is introduced. BUPTCMCCCMG-IMT2030 was developed based on the framework of the 5G GBSM standard channel model, and a series of extensions for the 6G technologies and scenarios described in Section IV-VII were integrated. The software structure, main functions, and a few simulation examples including configurations and results of the BUPTCMCCCMG-IMT2030 are given below.

### v4

@@ VI RIS Channel Measurement and Modeling
Different from traditional communication technologies, an RIS realizes active control of a wireless signal propagation environment by passively reflecting electromagnetic waves. Fig. 22 shows a typical RIS-assisted communication scenario. In this scenario, the LOS path is obstructed, and an RIS is employed to enhance the received signal. Obviously, traditional channel modeling methods are not suitable for this kind of channel, which requires new research on the RIS channel model. Notably, in Fig. 22, the RIS channel was characterized by a Tx-RIS-Rx concatenated channel, which was composed of a Tx-RIS subchannel, RIS-Rx subchannel, and RIS response. In this concatenated channel, the subchannels required independent modeling and were combined in the RIS response. The subsequent RIS channel modeling described in this paper is based on this method.
@@ VI-A RIS Channel Measurements
Several other measurement campaigns have been conducted in practical environments to investigate RIS-assisted channels. For example, near-field RIS channel measurements were performed in both indoor and outdoor settings to study the path loss characteristics of RIS channels [248]. To support system-level simulations, [249] proposes a physical model for reflective RIS and conducts corresponding indoor measurements. This work aims to abstract the RIS physical properties for integration into practical channel modeling frameworks. In [250], a RIS-assisted channel measurement platform based on time-domain correlation principles is developed, which can be applied to a variety of complex deployment scenarios. Utilizing this platform, the authors carry out a measurement campaign at $26$ GHz in a typical indoor corridor environment, as shown in Fig. 23. Based on the collected data, they analyze the path loss characteristics of RIS-assisted links under corridor conditions. Furthermore, the study in [251] investigates channel reciprocity in RIS-assisted communication networks under TDD systems. Through measurements conducted in indoor scenarios with two representative RIS types, the authors demonstrate that the channel reciprocity theorem generally holds. In addition, [252] and [253] conduct RIS channel measurements at $2.6$ GHz across indoor, outdoor, and O2I scenarios. The former analyzes the PDP, path loss, SF, and AS under different RIS configuration strategies, while the latter focuses on variations in the Rician $K$-factor introduced by RIS deployment.
@@
1 bit 5.8 Indoor, outdoor Power gain [248]
1 bit 5.8
@@
radiation pattern,
channel reciprocity
[247]
1 bit 26 Indoor PL [250]
1 bit 26.9 Indoor Radiation pattern [249]
Continuously 4.25, 10.5 Microwave anechoic chamber PL [254]
1 bit 27, 33 Microwave anechoic chamber
PL,
@@
Continuously, 1 bit 4.25, 27 Indoor Channel reciprocity [251]
1 bit 2.6 Indoor, outdoor, O2I PDP, PL, SF, DS, K [252, 253]
@@ VI-B RIS Channel New Characteristics
In the RIS channel, there are three types of characteristics that are worth paying attention to, namely the reflection coefficient characteristics of RIS element, RIS radiation pattern characteristics, and RIS concatenated loss characteristics. Among them, RIS radiation pattern characteristics consists of RIS overall radiation pattern characteristics and RIS element radiation pattern characteristics. The reasons are as follows.
Firstly, the RIS is one of the most important nodes in the RIS channel, and its response are an important component of the RIS channel. In order to calculate the RIS response, it is necessary to calculate the reflection coefficient of each RIS element in advance. Thus, the RIS element reflection coefficient characteristics are very important. Secondly, the final form of RIS response is RIS overall radiation pattern or RIS elemenmt radiation pattern, therefore, their characteristics can not be ignored either. Fig. 24 shows a feasible process of calculating the RIS radiation pattern, corresponding to (18)-(23). Finally, since the RIS concatenated channel is obtained by cascading Tx-RIS sub-channel, RIS response, and RIS-Rx sub-channel, its loss characteristics are different from traditional communication channel. Thus, RIS concatenated loss characteristics are worth studying.
@@ VI-B2 RIS radiation pattern characteristics
VI-B3 RIS concatenated loss characteristics
Multiple studies have been conducted on the RIS concatenated loss of the RIS channel based on theoretically derived and measurements. They provide rich RIS concatenated loss modeling methods and combined reveal the RIS concatenated loss characteristics of the RIS channel.
During a period, the relationship between the RIS concatenated loss of the RIS channel and the distances $d_{1}$ and $d_{2}$ is very relevant, where $d_{1}$ is the distance between the transmitter and RIS, and $d_{2}$ is the distance between the RIS and receiver. A theoretical RIS concatenated loss model in free space under the far-field condition was proposed in [261]. The theoretical analysis demonstrated that the RIS concatenated loss is proportional to $(d_{1}d_{2})^{2}$. Similar conclusions were also obtained in measurements of a corridor scene. Authors [250] carried out a channel measurement in a typical indoor corridor scenario and studied the RIS concatenated loss characteristic of the RIS channel in corridor scenarios. The results indicate that a modified version of the multiplicative distance model can effectively fit the measurement data.
Furthermore, some different conclusions can be made. In [246], a general RIS concatenated loss model and some special models in free space were proposed. These special models apply to RIS-assisted far-field beamforming, RIS-assisted near-field focusing, and RIS-assisted near-field broadcasting cases. The RIS-assisted far-field beamforming and RIS-assisted near-field broadcasting models were verified by several measurements conducted in a microwave chamber at 27 and 33 GHz. Different from above, the authors found that the RIS concatenated loss is proportional to $(d_{1}+d_{2})^{2}$ in the RIS-assisted near-field broadcasting case. This indicates that the RIS concatenated loss characteristics of the RIS channel are related to the encoding method of the RIS and depend on whether the terminal is in the far field or near field.
@@ VI-C RIS Channel Modeling
Considering the prevalent use of reflective RIS, it is often placed between the transmitter and receiver, resulting in a concatenated channel known as the Tx-RIS-Rx channel. The main objective of this subsection is to construct an effective concatenated RIS channel model to accurately capture the response characteristics of the RIS and channel properties, while balancing the model complexity and accuracy.
@@
In [245], the authors propose a GBSM-based channel modeling framework for the RIS concatenated channel, which takes into account the incident angle dependence of the RIS element reflection coefficient. The model was applied to an RIS-assisted MIMO communication scenario, as shown in Fig. 22. The CIR of the RIS concatenated channel between the $s^{\textrm{th}}$ transmitter and the $u^{\textrm{th}}$ receiver is expressed as (24). Similar to the channel model presented in (6), most of the channel parameters in the RIS concatenated channel model are defined similarly. However, the RIS concatenated channel model differs in that it includes two subchannels and the RIS response. The specific parameters of the RIS concatenated channel model are listed below.
@@
Implementation of the proposed model is based on the 3GPP standardized channel model with some improvements, including three parts: 1) general and large-scale parameters, 2) small-scale parameters, and 3) the CIR for RIS concatenated channel. The model generates large-scale parameters (LSPs) and SSPs for both the Tx-RIS and RIS-Rx subchannels. The overall radiation pattern of the RIS is calculated using a method proposed in [259] to account for the incident angle-dependent characteristics of the RIS reflection coefficient. For further details on the specific implementation processes, readers may refer to [245].
@@ VI-D Summary and Prospects
The RIS channel measurement is primarily aimed at measuring the power gain caused by RIS. The characteristics of the RIS channel primarily include reflection coefficient characteristics of the RIS element, RIS radiation pattern characteristics, and RIS concatenated loss characteristics. These characteristics are worth considering in RIS channel modeling. Based on these findings, we foresee some challenges that RIS channel modeling research could encounter.
@@ VI-D1 RIS channel measurement
RIS channel measurement plays a crucial role in understanding RIS channel characteristics and validating RIS channel models. However, due to the high transmission loss of the RIS-assisted channel caused by the Tx-RIS-Rx concatenated link, it imposes greater requirements on the transmission power and dynamic range of the measurement platform. In addition, the measurement platform needs a complete architecture, such as multiple channels, a high sampling rate, and transmit-receive separation, to capture channel characteristics in various complex scenarios.
@@ VI-D3 RIS channel modeling
The passive characteristic makes RIS have the advantage of low cost and low power consumption. But this also leads to a problem where RIS requires a large area to provide sufficient gain. Therefore, the RIS channel may exhibit near-field effect and SnS like XL-MIMO channel. In this case, the RIS response can not be characterized by the RIS overall radiation pattern. Thus, modeling near-field effects and SnS in the RIS channel is a major challenge.
@@ VII SAGIN Channel Measurements and Models
One of the 6G usage scenarios is ubiquitous connectivity by addressing the challenges of connectivity, coverage, capacity, data rate and the mobility of terminals [6]. One focus of this usage scenario is to address presently uncovered or scarcely covered areas, and SAGIN is considered a resolution to achieve worldwide coverage. Recently, the Working Party 4B of the ITU-R Study Group 4 (ITU-R SG4) held its 56th plenary session, where it discussed the time planning for satellite technology standards for 6G and determined the initial work plan [291]. SAGIN is an integration of satellite systems, aerial networks, and terrestrial communications. Thus in SAGIN, channel characteristics in the communication scenarios of different segments (space, air and ground) should be considered. However, due to the long propagation distance and high node mobility, SAGIN channel characteristics are distinctly different from those of the traditional terrestrial channels [292]. Thus in this section, the present research on channel measurements, characteristics, and models of SAGIN is reviewed. Since terrestrial systems have been discussed in the above sections, this section is focused more on the SG and AG scenarios. In addition, a brief summary and potential future research directions are provided.
@@ VII-A SAGIN Channel Measurements
For clarity, SAGIN channel measurements are classified into SG, AG, air-to-air (AA) and air-to-sea (AS), and summarized them in Table V. For SG channels, research work has been started in the early years. Earlier, researchers [266] conducted measurements using the ATS-6 satellite to study the propagation loss of SG links at 860 and 1550 MHz. In [267], propagation loss measurements were carried out at several geographical ground locations in the Q/V frequency band, and a low earth orbit (LEO) satellite channel model was developed using machine learning and statistical methods. A series of measurements of tree attenuation were conducted and studied in various frequency bands such as L-band and K-band [268, 269, 270, 271, 272, 273, 274]. In [275], a satellite MIMO channel study was conducted to investigate the channel capacity in an urban scenario. In [278], meteorological effects, such as rain, clouds, and scintillation, were studied based on field channel measurements in a tropical area. In addition, researchers in [276, 277] conducted measurements to study rain attenuation at the Ka/Q band. According to Table V, it can be found that most of the channel measurements of SG were carried out at L-band, S-band, K-band and Q-band [293], which are the common frequency bands for satellite communications. For channel characteristics, these measurement campaigns mainly focus on the path loss, shadow fading (especially shadow fading caused by tree attenuation), and meteorological effects, and less attention is paid to the small scale characteristics of the channel.
@@
For aerial channels like AG and AA, due to the rise of UAV technology, the corresponding channel measurements have gradually conducted. In [284], the path loss of the UAV channel was studied at 2.3 GHz in urban and nonurban scenarios. In [280, 281, 282], the National Aeronautics and Space Administration (NASA) Glenn Research Center sponsored a comprehensive AG channel measurement study of the C-bands. These measurements covered nearly all the typical local ground environments, including overwater, hilly, mountainous, suburban, and near-urban areas. Researchers in [283] also conducted wideband measurements using a direct-sequence spread-spectrum channel measurement platform to characterize propagation at 2.05 GHz. In [279], a channel measurement for a low-altitude AG channel was conducted at 915 MHz. In [285], AG channel measurements were conducted in open and suburban areas at 4.3 GHz. In [287], authors studied time-varying AG channels based on the measurement data at 6.5 GHz, with a particular focus on the clustering and tracking of multipath components. In [286], the measurement of the AG channel in a near-airport scenario at 5 GHz was conducted. In addition, AA channels for UAVs were investigated based on laboratory measurements and several experimental flights in [288]. [289] measured the short-range AS channel at 5 GHz in the Pacific Ocean using a channel sounder based on Direct Sequence-Spread Spectrum (DS-SS) technology, and verified the proposed non-stationary channel model for unmanned aerial vehicles to ships. [290] measured the AS channel at 2.2 GHz using NI USRP software-defined radio equipment and evaluated the communication performance of the system. According to Table V, it can be found that measurements of aerial channels are conducted mainly in AG scenarios, channel measurements in AA and other scenarios are scarce. Besides, measurement campaigns were conducted in relatively low frequency bands, which are from 0.92 to 6.5 GHz. For channel characteristics, in addition to path loss and shadow fading, AG channel measurements also focus on small scale characteristics such as delay spread, Ricean K-factor (K-factor), Doppler spread and others.
@@ VII-C SAGIN Channel Modeling
where $r$ is the received signal envelope, $ze^{j\phi_{0}}$ represents the LOS component, $we^{j\phi}$ represents other multipaths, the phases $\phi_{0}$ and $\phi$ are uniformly distributed between 0 and $2\pi$, $z$ is log-normally distributed , and $w$ has a Rayleigh distribution. In recent years, meteorological effects have been gradually taken into consideration. In [298], a channel model taking into account weather impairments was proposed. In this case, the PDF of the received signal amplitude is
@@
$K_{\rm total}=K_{\rm LOS}-K_{\rm Rain}-K_{\rm Cloud},$ (27)
@@ VII-D3 SAGIN channel modeling
Channel simulator software is used to generate CIRs in different scenarios and configurations on the basis of the existing modeling framework and configuration parameters. A channel simulator can play an important role in designing, evaluating, and optimizing communication systems. In this section, channel simulators capable of supporting potential 6G technologies are reviewed, and BUPTCMCCCMG-IMT2030 is introduced as an example of a 6G channel simulator. A series of extensions for the 6G technologies and scenarios described in Sections III–VII were integrated. The simulation principles, structure, functions and applications are given below.